# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v17

@@ 11.3 Possible Connections to Quantum Physics
For input distribution without bounding, Formula (2.3) may not converge, for example, when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$. Formula (2.3) converges in all cases only when the input uncertainty is bounded due to sampling, which resembles the need for re-normalization in quantum field theory [21]. The quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: what it is depends on how it is measured. For example, because mathematical divergence seems to correspond to distributional zero, it is even tempting to postulate if a localized experiment changes $\hat{\kappa}$ to produce distributional zero and the quantum result as “waveform collapse” [21]. Both resemblances are worth further investigation.
It is also worthwhile investigating if uncertainty bias produces Lamb shift [21].

### v18

@@ 11.3 Possible Connections to Quantum Physics
For input distribution without bounding, Formula (2.3) may not converge at all, for example, when the input uncertainty is Laplace [4], $\zeta(2n,\infty)=(2n)!$. Formula (2.3) converges in all cases only when the input uncertainty is bounded because of sampling, which resembles the need for re-normalization in quantum field theory [21]. Furthermore, the quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: What it is depends on how it is measured. For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse. Both resemblances are worthy of further investigation.